The sum of 3 number is 38. The second number is nine more than double the first, and the third number is one less than double the first.
Answer by stanbon(75887)

The sum of 3 number is 38. The second number is nine more than double the first, and the third number is one less than double the first.
------
Equations:
a + b + c = 38
b = 2a + 9
c = 2a-1
------
Substitute for "b" and for "c" and solve for "a":
a + (2a+9)+(2a-1) = 38
------
5a + 8 = 38
5a = 30
a = 6
-----
b = 2a+9 = 12+9 = 21
-----
c = 2a-1 = 12-1 = 11
--------------------------
Cheers,
Stan H.
==================